# Causal machine learning methods and use of sample splitting in settings with high-dimensional confounding

### v1

@@ 4.2 Targeted Maximum Likelihood Estimation
In contrast to AIPW, after initial estimation of the outcome and exposure models, TMLE has a targeting or updating step, again with the purpose of correcting the bias of the g-computation estimator (Zheng and van der Laan, 2011). Using a procedure already well described in the literature (e.g., Schuler and Rose, 2017), the fitted exposure score model is used to obtain targeted or updated predicted outcomes $\hat{E}^{*}_{1}(w)$ and $\hat{E}^{*}_{0}(w)$ for $X=1$ and $X=0$ respectively for each record, and these are used to estimate the ACE, with the estimator defined as
@@ 5.1 Data-adaptive approaches
Data-adaptive approaches can be used within the DR methods to estimate each of the nuisance functions, with the predictive performance of the approaches being the key criterion of interest. Therefore, nuisance function estimation can be viewed as a prediction modelling problem. One can consider parametric data-adaptive approaches, such as penalised regression methods (e.g., least absolute shrinkage and selection operator (LASSO) (Tibshirani, 1996; Hastie et al., 2001, pg. 64-65))) wherein coefficients can be shrunk towards zero (or in some cases, e.g., elastic net regression, set to zero), which are controlled or influenced by decision rules of the method. One can also consider non-parametric data-adaptive approaches, which include tree-based (e.g., random forest (RF) Breiman (2001)) and non-tree-based (e.g. neural networks (Kuhn and Johnson, 2013, pg. 141-145), support vector machines (Hastie et al., 2001, pg. 337-389)) methods, and often allow greater flexibility than the parametric approaches. In this work we consider both parametric and non-parametric data-adaptive approaches.
@@ 5.1.1 Ensemble Learning
In seeking to improve predictive performance, it is common to use an ensemble learning approach (Zhou, 2012, pg.15), wherein multiple data-adaptive approaches, termed learners, are considered simultaneously. The group of learners considered for the prediction problem is called a candidate library, and the learners within are used to solve the same prediction problem, but predictions from the learners are combined to arrive at a solution with strengthened predictive ability.
In this work, we use the Super Learner (SL), a widely used ensembling approach in the context of DR estimation of causal effects (Zivich and Breskin, 2021; Schuler and Rose, 2017; Lendle et al., 2013; Kreif and DiazOrdaz, 2019; Decruyenaere et al., 2020; Herrera et al., 2018; Meng and Huang, 2022).
@@ 5.2 Sample splitting
As mentioned previously, sample splitting has been proposed as an approach to help overcome issues that may arise from overfitting when using CML approaches (Kennedy, 2016; Zivich and Breskin, 2021). In this section, we outline a widely used form of sample splitting called cross-fitting (CF) (Chernozhukov et al., 2018; Kreif and DiazOrdaz, 2019; Jacob, 2020; Chernozhukov et al., 2020, 2017) that we consider in this work.
CF (or K-fold CF) is where the sample of size $n$ is split randomly into $K\geq 2$ parts (folds) of roughly equal size ($n/K$). The nuisance models are fitted on all but one fold (the complement), and the remaining fold is used to get predictions from these models. The process is repeated $K$ times, so that each fold $k=1,2,...,K$ is used to obtain predictions once (by rotation). The predictions are then used within the given DR method to obtain a final estimate of the causal effect (Chernozhukov et al., 2017). In this study, when CF was applied, standard errors were calculated from the influence curve (refer to Section 2 of the Supplementary material for implementation details).
@@ 6.1 Aim
We aimed to compare the performance of AIPW and TMLE using SL with different choices of library (reduced vs full) and different implementations of CF (no CF, and with 2, 5 and 10 folds) in realistic scenarios, largely informed by the BIS case study. We considered a range of scenarios, varying data generating mechanism complexity, confounder set size (p=14 or 87), and sample size (n=200, 500, 1000, 2000). The small confounder set was modelled on demographic and background confounders in the BIS study, and the large set additionally considered metabolites (see Section 3 of the Supplementary material (Tables S2-S4) for more details).
@@ 6.2 Data-generating mechanisms
The variables (confounders, exposure, and outcome) were generated in the order specified in Section 4 of the Supplementary material (Tables S5-S7). Parametric regression models were used to simulate variables for all five data-generating mechanisms, three of which involved the small confounder set (simple-1, complex-1a and complex-1b), and two of which involved the large confounder set (simple-2, complex-2), detailed below. Parameters for the data-generation models were, unless otherwise stated, obtained by fitting analogous models to the BIS data.
Under the simple data-generating mechanisms (simple-1 and simple-2), we considered exposure and outcome regression models with main effects only and a linear association for the continuous covariates. Complex data-generating mechanisms considered exposure and outcome models with main effects and two-way interactions (confounder-confounder and exposure-confounder as appropriate), with linear and non-linear associations for the continuous covariates. For the small confounder set, we used two complex scenarios, with coefficient values for interaction terms two times larger than observed in the BIS data (complex-1a) and another with coefficients for the interaction terms four times larger (complex-1b) than observed in the BIS data. For the large confounder set, we considered one complex scenario with coefficient values as observed in the BIS data (complex-2). For full details regarding the data-generating mechanisms refer to Section 4 of the Supplementary material (Tables S6-S11).
We simulated 2000 datasets for each data generation mechanism and sample size, with this number of simulation replications determined based on a requirement that the Monte Carlo SE for a coverage of 95% be no greater than 0.5%. Overall, 20 scenarios were considered (four sample sizes for five data-generating mechanisms). When generating the outcome in each scenario, we set the main effect to be of a size determined such that the null hypothesis of zero average causal effect was formally rejected (using the threshold of $p<0.05$) in approximately 80% of the simulated datasets for the given sample size. The intercepts in all outcome models were modified so that the mean of $Y$ remained at 0. Further details on the parameters used in the generation of confounders, exposure, and outcome are provided in Section 4 of the Supplementary material (Tables S12-S23).
@@ 6.3 Estimand and methods compared
We estimated the average causal effect (ACE) of X on Y via AIPW and TMLE with no CF and with CF, using varying folds ($2$, $5$ and $10$). SL was used with 10-fold cross-validation for outcome and exposure prediction modelling. Reduced and full candidate libraries for SL were considered using default hyperparameters for each learner. The reduced library included non-adaptive and data-adaptive parametric approaches. The full library included approaches in the reduced library as well as non-parametric data-adaptive approaches. Further details on the SL candidate libraries are provided in Section 5 of the Supplementary material (Table S24). To deal with extreme inverse probability weights, truncation of propensity score predictions at the 5th and 95th percentiles was undertaken.
@@ 6.4 Performance measures
For each scenario, to evaluate performance, we calculated the bias (the difference between the average of the ACE estimates across the 2000 simulations and the true value of the ACE), the relative bias (bias divided by the true ACE, as percent) as well as the empirical standard error, average model-based standard error (model SE) and relative error of the model-based SE compared to the empirical SE (%). We also estimated the coverage probability of the 95% Wald CI. To calculate performance measures, the true values of the ACE used in each scenario were obtained empirically in a single very large simulated dataset (full details and true values used are provided in Section 6 of the Supplementary material (Table S25)). We also estimated the Monte Carlo Standard Errors (MCSE’s) for each performance measure. Morris et al. (2019) provide more detail on the performance measures and how they are obtained.
@@ 7 Simulation study: results
Our presentation and assessment of performance focus on the reporting of bias in point estimates (relative bias), the empirical SE, relative error in the model-based SE, and coverage probability. In general, patterns for the complex-1b scenario were similar to complex-1a. Hence, Figures LABEL:fig:Relbias_red-LABEL:fig:CP_red show performance measures for AIPW and TMLE when the reduced and full libraries were applied in SL across sample sizes, for the small confounder set (simple-1 and complex-1a) and the large confounder set (simple-2 and complex-2). Tables S26-S34 in Section 7 of the Supplementary material contain results for both the reduced and full libraries across all scenarios, for the comprehensive panel of performance measures, including associated Monte Carlo standard errors. This section mainly describes the results, and a fuller discussion of the findings is provided in Section 9.
@@
In general, both DR methods performed similarly across the key performance measures, except that AIPW failed to produce sensible results for a few datasets when CF was applied (2-12 datasets across scenarios). In these settings, the standard error was $>10$ times the median standard error or the absolute value of the point estimate was $>5$ times the absolute value of the median point estimate. These results were omitted from the figures, with details provided in Figures LABEL:fig:Relbias_red-LABEL:fig:CP_red and in the Supplementary material (Section 7, Table S26).
@@ 7.2 Impact of using cross-fitting
In general, relative bias appeared to be of similar magnitude whether CF was applied or not (Figure LABEL:fig:Relbias_red). Empirical SE was observed to be similar with or without CF, except for complex-2 at smaller sample sizes (n=200, 500), where it was larger with CF than without (Figure LABEL:fig:EmpSE_red). In general, the use of CF resulted in lower relative error in model-SE than without CF (Figure LABEL:fig:RelmodSE_red). The use of CF appeared to improve coverage compared to not using CF, resulting in coverage probabilities that were closer to nominal (Figure LABEL:fig:CP_red).
@@ 7.3 Impact of varying the number of folds within cross-fitting
In general, method performance was similar regardless of the folds used within CF (Figures LABEL:fig:Relbias_red-LABEL:fig:CP_red). An exception was noted at larger sample sizes (e.g., n=2000) with complex-1a, where using 2 folds appeared to result in lower bias compared to the use of $\geq$ 5 folds with CF (Figure 2). Across scenarios at smaller sample sizes, empirical SE was slightly smaller when using $\geq$ 5 folds than with 2 folds, but at larger sample sizes (e.g., n=2000), empirical SE was similar regardless of the number of folds used (Figure 3). Patterns were harder to discern for the relative error in model-SE (Figure 4), with other factors appearing more important than the number of folds (e.g., the method). For coverage, with the larger confounder set (simple-2 and complex-2), fewer folds appeared slightly better as sample size increased (Figure 5).
@@ 8 Application to the BIS case study
For the BIS case study example, we applied both AIPW and TMLE without CF and with CF, using 2, 5 and 10 folds. We used reduced and full libraries in SL (composition detailed in Section 5 of the Supplementary material (Table S24), and we considered the two confounder sets analogous with those that motivated the simulation study design (detailed in Section 3 of the Supplementary material (Tables S2-S4)). Henceforth, we use the following terminology: partially adjusted refers to estimates obtained using a confounder set that contains demographic and background confounders only (p=14) and fully adjusted refers to an estimate adjusted for a set that contains demographic, background, and metabolomic confounders (p=87).
@@ 8.1 Results
The results are presented in Figure LABEL:Fig_res. Full results are available in Section 8 of the Supplementary Material (Table S35, and the distribution of estimated propensity scores by exposure group are provided in Figures S1 and S2).
@@ 9 Discussion
We conducted a simulation study that was motivated by a realistic study to answer key questions regarding the relative performance of AIPW and TMLE, the use of CF, and the number of folds used within CF, when estimating the ACE in the presence of high-dimensional confounding. We found that in general in this setting, AIPW and TMLE performed similarly but AIPW exhibited some stability issues. CF was not that important for the bias of point estimates but was important to reduce the bias in the model-based SE, and for improving the coverage of the methods. In general, there was no substantial difference between using 2, 5 or 10 folds when using CF with the CML methods. In complex scenarios, the full library was important for bias in point estimates, especially at larger sample sizes, and for empirical SE.
@@
In the high-dimensional setting our findings suggest that, although either AIPW or TMLE may be sensible choices of method to estimate the ACE, some caution is warranted with AIPW due to some of its vulnerabilities. This is likely explained by limitations of AIPW, where estimates produced may not always be within the natural bounds of the problem, in contrast to TMLE where this is not an issue because it is a substitution estimator (Díaz, 2019). With a complex scenario (regardless of the number of confounders), the full library appeared important, and this is as we would expect, with the more flexible algorithms needed to capture the complexity. Regardless of sample size or the number of confounders used, CF should be used with CML methods, and this is even more important if a full library is used. These results are in line with expectations that CF helps reduce the bias in SE estimation arising from the use of flexible ML algorithms. Interestingly, the number of folds used in CF did not have much impact on method performance, with other factors appearing more important (e.g. method and library choice).
@@
Our findings align with previous methodological studies evaluating and comparing these CML methods with and without CF, although, unlike our study, most have considered low-dimensional confounding only (with very few or binary confounders only), with little focus on strong high-dimensional confounding (Naimi et al., 2023; Zivich and Breskin, 2021) or motivation from realistic case studies. One exception is a recent study by Meng and Huang (2022) which considered both a realistic setting and high-dimensional confounding ($\approx$1000 observations, 331 covariates), and like us, observed that AIPW and TMLE performed similarly in general. Li et al. (2022) evaluated CML methods, finding that both AIPW and TMLE with CF are favourable for robust inferences as we found, but their study was in the context of a randomised experiment, with sample sizes in general much larger than we considered
@@
Meng and Huang (2022) found that parametric learners generally performed similarly to or outperformed non-parametric learners in a realistic setting and had lower computational times. Our findings in simple scenarios are consistent with that study, but in complex scenarios (e.g., complex-2), we found that the use of the full library was important to reduce bias in point estimates and variance, particularly for larger sample sizes, with variation depending on number of confounders used. An explanation for this could be that our full library considered learners that extended beyond tree-based methods, with the importance of doing so highlighted in a recent commentary (Balzer and Westling, 2023). It could also be due to differences in the design of our simulation study such as our data generating mechanisms, because conclusions in regard to the use of parametric versus non-parametric learners will likely be very specific to the chosen data-generating mechanisms, or due to the variant of CF we used (single CF) as opposed to other variants of CF available including what is referred to as double CF (Zivich and Breskin, 2021). Phillips et al. (2022) provide guidance regarding the choices that need to be made when using SL, including library construction. They state the importance of using a diverse library in SL and that inclusion of additional learners is not generally problematic. However, they emphasise that the number of learners included in a library might need to be limited in the high-dimensional confounding setting, or rather when the effective sample size, the ratio between sample size and the number of confounders, is low. Our study adds to this important literature, by re-emphasising the need to carefully consider the setting specific constraints (e.g., sample size, number of confounders) when choosing a library.
In the literature it has been suggested that the number of folds considered with single CF should be higher for smaller sample sizes than for larger sample sizes (Balzer and Westling, 2023). Until our study, there has been no existing guidance on the number of folds to use in CF in realistic settings that feature high-dimensional confounding (although in simple settings without high-dimensional confounding for example, Chernozhukov et al. (2018) report that 4 or 5 may work better than 2). Our study found that in general, CML method performance was similar regardless of the number of folds used in CF. Even though for coverage, fewer folds appeared slightly better as sample size increased with more confounders, in line with previous recommendations, other factors (library, method) appeared more important across the full range of performance indicators.
@@
Our research strongly suggests that caution is warranted when applying these methods, with careful consideration of SL library composition required. We found indications that performance does not necessarily improve with increasing sample sizes. For example, in scenarios where we considered the large confounder set and used CF, we found increasing bias in model-SE and decreasing coverage, alongside decreasing bias in point estimates as larger sample sizes were considered. For these scenarios (simple-2, complex-2), we noted larger SL coefficients were assigned to the glmnet learner at smaller sample sizes than at larger samples sizes (Supplementary Section 9, Figures S3 and S4). Here we also observed that the estimated propensity scores were more variable and had a much smaller minimum (closer to zero) for the larger sample sizes than for the smaller sample sizes (Supplementary Section 9, Figure S5 and S6). A possible explanation for this observation is that at the lower sample size, the screening within glmnet reduces the variable set, inducing bias in point estimates but dampening variability in the propensity scores. At larger sample sizes the retention of variables helps reduce bias in point estimates, but induces greater variability, with estimated propensity scores closer to zero resulting in biased model-SE’s and poorer coverage. This phenomenon could improve when considering sample sizes much larger than we considered. However, when faced with limited sample sizes in modern observational studies, it can be postulated that improved results could be obtained by carefully examining and selecting the learners within SL. For example, one could consider including ridge regression or high-dimensional regression (Sur and Candès, 2019) as learners, as they are better suited to handling a large number of (“non-null") variables.

### v2

@@ 4.2 Targeted Maximum Likelihood Estimation
In contrast to AIPW, after initial estimation of the outcome and exposure models, TMLE has a targeting or updating step, again with the purpose of correcting the bias of the g-computation estimator. 24 Using a procedure already well described in the literature (e.g., Schuler and Rose2, Luque-Fernandez et al 20), the fitted exposure score model is used to obtain targeted or updated predicted outcomes $\hat{E}^{*}_{1}(w)$ and $\hat{E}^{*}_{0}(w)$ for $X=1$ and $X=0$ respectively for each record, and these are used to estimate the ACE, with the estimator defined as
@@ 5.1 Data-adaptive approaches
Data-adaptive approaches can be used within the DR methods to estimate each of the nuisance functions, with the predictive performance of the approaches being the key criterion of interest. Therefore, nuisance function estimation can be viewed as a prediction modelling problem. One can consider parametric data-adaptive approaches, such as penalised regression methods (e.g., least absolute shrinkage and selection operator (LASSO) 43 44 (p 64-65)) wherein coefficients can be shrunk towards zero (or in some cases, e.g., elastic net regression, set to zero), which are controlled or influenced by decision rules of the method. One can also consider non-parametric data-adaptive approaches, which include tree-based (e.g., random forest (RF) 45) and non-tree-based (e.g. neural networks 46 (p 141-145), support vector machines 44 (p 337-389)) methods, and often allow greater flexibility than the parametric approaches. In this work we consider both parametric and non-parametric data-adaptive approaches.
@@ 5.1.1 Ensemble Learning
In seeking to improve predictive performance, it is common to use an ensemble learning approach 47 (p 15), wherein multiple data-adaptive approaches, termed learners, are considered simultaneously. The group of learners considered for the prediction problem is called a candidate library, and the learners within are used to solve the same prediction problem, but predictions from the learners are combined to arrive at a solution with strengthened predictive ability.
In this work, we use the Super Learner (SL), a widely used ensembling approach in the context of DR estimation of causal effects. 3, 2, 18, 27, 28, 29, 26
@@ 5.2 Sample splitting
As mentioned previously, sample splitting has been proposed as an approach to help overcome issues that may arise from overfitting when using CML approaches. 22, 3 In this section, we outline a widely used form of sample splitting called cross-fitting (CF) 23, 27, 48, 49, 11 that we consider in this work.
CF (or K-fold CF) is where the sample of size $n$ is split randomly into $K\geq 2$ parts (folds) of roughly equal size ($n/K$). The nuisance models are fitted on all but one fold (the complement), and the remaining fold is used to get predictions from these models. The process is repeated $K$ times, so that each fold $k=1,2,...,K$ is used to obtain predictions once (by rotation). The predictions are then used within the given DR method to obtain a final estimate of the causal effect. In this study, when CF was applied, standard errors were calculated from the influence curve (refer to Section 2 of the Supporting Information for implementation details). 11 For TMLE with CF, targeting was performed in the remaining data that was not used to fit the nuisance models, an approach consistent with van der Laan and Rose 9, Motoya et al50 and Balzer and Westling 12, although we note that there are alternative implementations which retain the same theoretical properties. 51
@@ 6.1 Aim
We aimed to compare the performance of AIPW and TMLE using SL with different choices of library (reduced vs full) and different implementations of CF (no CF, and with 2, 5 and 10 folds) in realistic scenarios, largely informed by the BIS case study. We considered a range of scenarios, varying data generating mechanism complexity, confounder set size (p=14 or 87), and sample size (n=200, 500, 1000, 2000). The small confounder set was modelled on demographic and background confounders in the BIS study, and the large set additionally considered metabolites (see Section 3 of the Supporting Information (Tables S2-S4) for more details).
@@ 6.2 Data-generating mechanisms
The variables (confounders, exposure, and outcome) were generated in the order specified in Section 4 of the Supporting Information (Tables S5-S7). Parametric regression models were used to simulate variables for all five data-generating mechanisms, three of which involved the small confounder set (simple-1, complex-1a and complex-1b), and two of which involved the large confounder set (simple-2, complex-2), detailed below. Parameters for the data-generation models were, unless otherwise stated, obtained by fitting analogous models to the BIS data.
Under the simple data-generating mechanisms (simple-1 and simple-2), we considered exposure and outcome regression models with main effects only and a linear association for the continuous covariates. Complex data-generating mechanisms considered exposure and outcome models with main effects and two-way interactions (confounder-confounder and exposure-confounder as appropriate), with linear and non-linear associations for the continuous covariates. For the small confounder set, we used two complex scenarios, with coefficient values for interaction terms two times larger than observed in the BIS data (complex-1a) and another with coefficients for the interaction terms four times larger (complex-1b) than observed in the BIS data. For the large confounder set, we considered one complex scenario with coefficient values as observed in the BIS data (complex-2). For full details regarding the data-generating mechanisms refer to Section 4 of the Supporting Information (Tables S6-S11).
We simulated 2000 datasets for each data generation mechanism and sample size, with this number of simulation replications determined based on a requirement that the Monte Carlo SE for a coverage of 95% be no greater than 0.5%. Overall, 20 scenarios were considered (four sample sizes for five data-generating mechanisms). When generating the outcome in each scenario, we set the main effect to be of a size determined such that the null hypothesis of zero average causal effect was formally rejected (using the threshold of $p<0.05$) in approximately 80% of the simulated datasets for the given sample size. We chose to keep the power constant to ensure that coverage probabilities, our main metric of interest, were comparable across sample size scenarios. The intercepts in all outcome models were modified so that the mean of $Y$ remained at 0. Further details on the parameters used in the generation of confounders, exposure, and outcome are provided in Section 4 of the Supporting Information (Tables S12-S23).
@@ 6.3 Estimand and methods compared
We estimated the average causal effect (ACE) of X on Y via AIPW and TMLE with no CF and with CF, using varying folds ($2$, $5$ and $10$). SL was used with 10-fold cross-validation for outcome and exposure prediction modelling. Reduced (less flexible) and full (more flexible) candidate libraries for SL were considered using default hyperparameters for each learner. The reduced library was chosen to include less flexible approaches. Specifically, it included the following non-adaptive and data-adaptive parametric approaches: GLM, GLM with pairwise interactions, Bayesian GLM, generalised additive modelling and GLM with Lasso/Elastic net regularisation. The full library was comprised of the approaches in the reduced library and an additional set of more flexible, non-parametric data-adaptive approaches, including learners more likely to violate the convergence rate and Donsker conditions required for obtaining valid inferences without cross-fitting. Further details on the SL candidate libraries are provided in Section 5 of the Supporting Information (Table S24). To deal with extreme inverse probability weights, truncation of propensity score predictions at the 5th and 95th percentiles was undertaken.
@@ 6.4 Performance measures
For each scenario, to evaluate performance, we calculated the bias (the difference between the average of the ACE estimates across the 2000 simulations and the true value of the ACE), the relative bias (bias divided by the true ACE, as percent) as well as the empirical standard error, average model-based standard error (model SE) and relative error of the model-based SE compared to the empirical SE (%). We also estimated the coverage probability of the 95% Wald CI. To calculate performance measures, the true values of the ACE used in each scenario were obtained empirically in a single very large simulated dataset (full details and true values used are provided in Section 6 of the Supporting Information (Table S25)). We also estimated the Monte Carlo Standard Errors (MCSE’s) for each performance measure. Morris et al 52 provide more detail on the performance measures and how they are obtained.
@@ 7 Simulation study: results
Our presentation and assessment of performance focus on the reporting of bias in point estimates (relative bias), the empirical SE, relative error in the model-based SE, and coverage probability. In general, patterns for the complex-1b scenario were similar to complex-1a. Hence, Figures 2-5 show performance measures for AIPW and TMLE when the reduced and full libraries were applied in SL across sample sizes, for the small confounder set (simple-1 and complex-1a) and the large confounder set (simple-2 and complex-2). Tables S26-S34 in Section 7 of the Supporting Information contain results for both the reduced and full libraries across all scenarios, for the comprehensive panel of performance measures, including associated Monte Carlo standard errors. This section mainly describes the results, and a fuller discussion of the findings is provided in Section 9.
Figure 2: Simulation study results for the relative bias of point estimates (%) for AIPW and TMLE with varying use of cross-fitting, by scenario (data generating mechanism and sample size). Section 5 of the Supporting Information (Table S25) provides the true value of the ACE used for each scenario. Note: results from a small number of datasets were excluded for AIPW (but not TMLE) (refer to Section 7.1, and Section 5 of the Supporting Information for more details).
Figure 3: Simulation study results for the Empirical SE for AIPW and TMLE with varying use of cross-fitting, by scenario (data generating mechanism and sample size). Note: results from a small number of datasets were excluded for AIPW (but not TMLE) due to the large effect that they had on some calculated performance measures (refer to Section 7.1, and Section 5 of the Supporting Information for more details).
Figure 4: Simulation study results for the bias in model-based SE (%) for AIPW and TMLE with varying use of cross-fitting, by scenario (data generating mechanism and sample size). Note: results from a small number of datasets were excluded for AIPW (but not TMLE) due to the large effect that they had on some calculated performance measures (refer to Section 7.1, and Section 5 of the Supporting Information for more details).
Figure 5: Simulation study results for the coverage probability (%) for AIPW and TMLE with varying use of cross-fitting, by scenario (data generating mechanism and sample size). Note: results from a small number of datasets were excluded for AIPW (but not TMLE) due to the large effect that they had on some calculated performance measures (refer to Section 7.1, and Section 5 of the Supporting Information for more details).
@@
In general, both DR methods performed similarly across the key performance measures, except that AIPW failed to produce sensible results for a few datasets when CF was applied (2-12 datasets across scenarios). In these settings, the standard error was $>10$ times the median standard error or the absolute value of the point estimate was $>5$ times the absolute value of the median point estimate. These results were omitted from the figures, with details provided in Figures 2-5 and in the Supporting Information (Section 7, Table S26).
@@ 7.2 Impact of using cross-fitting
In general, relative bias appeared to be of similar magnitude whether CF was applied or not (Figure 2). Empirical SE was observed to be similar with or without CF, except for complex-2 at smaller sample sizes (n=200, 500), where it was larger with CF than without (Figure 3). In general, the use of CF resulted in lower relative error in model-SE than without CF (Figure 4). The use of CF appeared to improve coverage compared to not using CF, resulting in coverage probabilities that were closer to nominal (Figure 5).
@@ 7.3 Impact of varying the number of folds within cross-fitting
In general, method performance was similar regardless of the folds used within CF (Figures 2-5). An exception was noted at larger sample sizes (e.g., n=2000) with complex-1a, where using 2 folds appeared to result in lower bias compared to the use of $\geq$ 5 folds with CF (Figure 2). Across scenarios at smaller sample sizes, empirical SE was slightly smaller when using $\geq$ 5 folds than with 2 folds, but at larger sample sizes (e.g., n=2000), empirical SE was similar regardless of the number of folds used (Figure 3). Patterns were harder to discern for the relative error in model-SE (Figure 4), with other factors appearing more important than the number of folds (e.g., the method). For coverage, with the larger confounder set (simple-2 and complex-2), fewer folds appeared slightly better as sample size increased (Figure 5).
@@ 8 Application to the BIS case study
For the BIS case study example, we applied both AIPW and TMLE without CF and with CF, using 2, 5 and 10 folds. We used reduced and full libraries in SL (composition detailed in Section 5 of the Supporting Information (Table S24), and we considered the two confounder sets analogous with those that motivated the simulation study design (detailed in Section 3 of the Supporting Information (Tables S2-S4)). Henceforth, we use the following terminology: partially adjusted refers to estimates obtained using a confounder set that contains demographic and background confounders only (p=14) and fully adjusted refers to an estimate adjusted for a set that contains demographic, background, and metabolomic confounders (p=87).
@@ 8.1 Results
The results are presented in Figure 6. Full results are available in Section 8 of the Supporting Information (Table S35, and the distribution of estimated propensity scores by exposure group are provided in Figures S1 and S2).
Figure 6: Estimated average causal effect (ACE) with accompanying $95\%$ CI, of inflammation (GlycA) in 1-year old infants on Pulse Wave velocity (PWV) at 4 years of age (standardised), obtained by applying the methods to the BIS motivating example. $\hat{}$ demographic and background confounders only; demographic, background, and metabolomic confounders.
Note: Estimate for AIPW using fully adjusted confounder set and full library in SL with 10 fold cross-fitting not shown in figure as too large to be sensible (refer to Table 35 in Section 8 of the Supporting Information).
@@ 9 Discussion
We conducted a simulation study that was motivated by a realistic study to answer key questions regarding the relative performance of AIPW and TMLE, the use of CF, and the number of folds used within CF, when estimating the ACE in the presence of high-dimensional confounding. We found that in general in this setting, AIPW and TMLE performed similarly but AIPW exhibited some stability issues. CF was not that important for the bias of point estimates but was important to reduce the bias in the model-based SE, and for improving the coverage of the methods. In general, there was no substantial difference between using 2, 5 or 10 folds when using CF with the CML methods. In complex scenarios, the full, more flexible library was important for bias in point estimates, especially at larger sample sizes, and for empirical SE.
@@
In the high-dimensional setting our findings suggest that, although either AIPW or TMLE may be sensible choices of method to estimate the ACE, some caution is warranted with AIPW due to some of its vulnerabilities. This is likely explained by limitations of AIPW, where estimates produced may not always be within the natural bounds of the problem, in contrast to TMLE where this is not an issue because it is a substitution estimator.6 With a complex scenario (regardless of the number of confounders), the full library appeared important, and this is as we would expect, with the more flexible algorithms needed to capture the complexity. Regardless of sample size or the number of confounders used, CF should be used with CML methods, and this is even more important if a full library is used. These results are in line with expectations that CF helps reduce the bias in SE estimation arising from the use of flexible ML algorithms. Interestingly, the number of folds used in CF did not have much impact on method performance, with other factors appearing more important (e.g. method and library choice).
@@
Our findings align with previous methodological studies evaluating and comparing these CML methods with and without CF, although, unlike our study, most have considered low-dimensional confounding only (with very few or binary confounders only), with little focus on strong high-dimensional confounding 1, 3 or motivation from realistic case studies. One exception is a recent study by Meng and Huang26 which considered both a realistic setting and high-dimensional confounding ($\approx$1000 observations, 331 covariates), and like us, observed that AIPW and TMLE performed similarly in general. Li et al53 evaluated CML methods, finding that both AIPW and TMLE with CF are favourable for robust inferences as we found, but their study was in the context of a randomised experiment, with sample sizes in general much larger than we considered. In the literature it has been suggested that the number of folds considered with single CF should be higher for smaller sample sizes than for larger sample sizes. 12 Until our study, there has been no existing guidance on the number of folds to use in CF in realistic settings that feature high-dimensional confounding (although in simple settings without high-dimensional confounding for example, Chernozhukov et al 23 report that 4 or 5 may work better than 2). Our study found that in general, CML method performance was similar regardless of the number of folds. Although fewer folds appeared to produce slightly better coverage as sample size increased with more confounders, in line with previous recommendations, other factors (library, method) appeared more important across the full range of performance indicators.
@@
Meng and Huang 26 found that parametric learners generally performed similarly to or outperformed non-parametric learners in a realistic setting and had lower computational times. Our findings in simple scenarios are consistent with that study, but in complex scenarios (e.g., complex-2), we found that the use of the full, more flexible library was important to reduce bias in point estimates and variance, particularly for larger sample sizes, with variation depending on number of confounders used. An explanation for this could be that our full library considered learners that extended beyond tree-based methods, with the importance of doing so highlighted in a recent commentary. 12 It could also be due to differences in the design of our simulation study such as our data generating mechanisms, because conclusions in regard to the use of parametric versus non-parametric learners will likely be very specific to the chosen data-generating mechanisms, or due to the variant of CF we used (single CF) as opposed to other variants of CF available including what is referred to as double CF. 3 Phillips et al 54 provide guidance regarding the choices that need to be made when using SL, including library construction. They state the importance of using a diverse library in SL though emphasise that the number of learners included might need to be limited in the high-dimensional confounding setting. Our study adds to this important literature, by re-emphasising the need to carefully consider the setting specific constraints (e.g., sample size, number of confounders) when choosing a library.
@@
Our research strongly suggests that caution is warranted when applying these methods, with careful consideration of SL library composition required. We found indications that performance does not necessarily improve with increasing sample sizes. For example, in scenarios where we considered the large confounder set and used CF, we found increasing bias in model-SE and decreasing coverage, alongside decreasing bias in point estimates as larger sample sizes were considered. For these scenarios (simple-2, complex-2), we noted larger SL coefficients were assigned to the glmnet learner at smaller sample sizes than at larger samples sizes (Supplementary Section 9, Figures S3 and S4). Here we also observed that the estimated propensity scores were more variable and had a much smaller minimum (closer to zero) for the larger sample sizes than for the smaller sample sizes (Supplementary Section 9, Figure S5 and S6). A possible explanation for this observation is that at the lower sample size, the screening within glmnet reduces the variable set, inducing bias in point estimates but dampening variability in the propensity scores. At larger sample sizes the retention of variables helps reduce bias in point estimates, but induces greater variability, with estimated propensity scores closer to zero resulting in biased model-SE’s and poorer coverage. This phenomenon suggests that careful selection of the learners within SL is important in the sample sizes commonly encountered in modern observational studies. Future research should examine inclusion of ridge regression or high-dimensional regression 55 as learners, which are expected to be better at handling a large number of variables with non-null effects.